Let R be a relation over set A, then: Prove that IA∘R=R=R∘IA. 


1
Expert's answer
2021-04-29T17:10:42-0400

As R is a relation on set A.


It implies that The elemnts of A must belong to the R in cartesian product form.


Taking LHS

"IA*R=A*R=R"


Since The product of the set on which relation is defined is equal to that relation.


Now taking LHS-

"R*IA=R*A=R"


This implies that-

"IA*R=R=R*IA" , Hence proved.